OpenStudy (patrickweewoos):

Which equations could model this problem? A package of 12 cans of dog food costs $33. What is the unit price d per can of dog food? Choose all answers that are correct.

OpenStudy (patrickweewoos):

A.)12d=33 B.)b divided by 12=33 C.)33/d=12 D.)33d=12

OpenStudy (igreen):

You have to divide..so you can use the equations: \(\dfrac{12}{33} = d\) or \(33d = 12\)

OpenStudy (patrickweewoos):

Ok @iGreen Its multiple answer question so its D and

OpenStudy (igreen):

I think it's only D..all the rest look incorrect.

OpenStudy (igreen):

Is option B this? 12/d = 33

OpenStudy (igreen):

Or is it d/12 = 33?

OpenStudy (patrickweewoos):

d divided by 12 =33

OpenStudy (igreen):

Okay, that's wrong..so D is the only option..

OpenStudy (patrickweewoos):

Six friends go out to lunch and decide to split the bill evenly. The bill comes to $61.56. How much does each person owe?

OpenStudy (patrickweewoos):

@iGreen

OpenStudy (igreen):

Just divide. 61.56 / 6 = ?

OpenStudy (patrickweewoos):

10.26?

OpenStudy (igreen):

Yep!